A ball is hit so that it travels straight upward after being struck by the bat. A fan observes that it takes 3.00 s for the ball to reach its maximum height. Find (a) the ball’s initial velocity and (b) the height it reaches.
1
Expert's answer
2020-10-09T07:07:16-0400

solution

given data

time taken to achieve maximum height(t)=3 sec

by the first equation of motion

"v=u+gt"

at maximum height v=0 m/s

g= - 9.8 m/s^2


putt the value


"u=9.8\times 3\\u=29.4m/s"


by the second equation of motion


"H_{max}=ut+\frac{1}{2}gt^2"


"H_{max}=29.4\times3-\frac{1}{2}\times 9.8\times3^2\\=44.1m"

so initial velocity is 29.4 m/s and maximum height is 44.1 m .